# Retrieving the hot CGM physics from the X-ray radial profile from eROSITA with an IlustrisTNG-based forward model

### v1

@@ 3.2 The hot gas component
In this work, we model the hot gas emission using the TNG300 hydrodynamical simulations (Pillepich et al. 2018; Marinacci et al. 2018; Naiman et al. 2018; Nelson et al. 2015; Springel et al. 2018); we use TNG300 to construct a lightcone and generate mock X-ray observations, as presented in Shreeram et al. (2024). Here, we summarize the most important features. We use the IllustrisTNG cosmological hydrodynamical simulation with the box of side length $302.6$ Mpc (Nelson et al. 2019, TNG300;) [^1] http://www.tng-project.org ; this box size allows us to map the hot CGM around MW-mass analogues embedded in the large scale structure. TNG300 contains $2500^{3}$ dark matter particles, with a baryonic mass resolution of $1.1\times 10^{7}\rm\ M_{\odot}$ (resulting in $\gtrsim 10^{3}$ particles at MW-mass galaxies), a comoving value of the adaptive gas gravitational softening length for gas cells of $370$ comoving parsec (allowing us to resolve the X-ray gas from $\sim 5$ kpc from the halo centre), gravitational softening of the collisionless component of $1.48$ kpc, and dark matter mass resolution of $5.9\times 10^{7}\ \rm M_{\odot}$. The TNG simulations adopt the Planck Collaboration XIII 2016 cosmological parameters. The TNG300 lightcone, LC-TNG300, is constructed with the box remap technique (Carlson & White 2010), and spans across redshifts $0.03\lesssim z\lesssim 0.3$; this range is motivated by observations (e.g. Comparat et al. 2022; Chadayammuri et al. 2022; Zhang et al. 2024a). It goes out to $1231\rm\ cMpc$ along the x-axis, subtending an area of $47.28\rm\ deg^{2}$ on the sky in the y-z plane. The physical properties of the distinct halos and subhalos within the TNG300 lightcone are obtained by the subfind algorithm (Springel et al. 2001; Dolag et al. 2009). subfind detects gravitationally bound substructures, equivalent to galaxies in observations and also provides us with a classification of subhalos into centrals and satellites, where centrals are the most massive substructure within a distinct halo. For the MW-mass bin [^2] This paper defines the stellar mass used from TNG300 as the mass within twice the stellar half-mass radius., $\rm M_{\star}=10^{10.5-11}{\rm M}_{\odot}$, we have $5,109$ centrals and $2,719$ satellites, resulting in a total simulated galaxy catalog with $7,828$ galaxies (subhalos).
@@
The X-ray photons are simulated within the LC-TNG300 in the $0.5-2.0$ keV intrinsic band with pyXsim (ZuHone & Hallman 2016), which is based on phox (Biffi et al. 2013, 2018), by assuming an input emission model where the hot X-ray emitting gas is in collisional ionization equilibrium. The spectral model computations of hot plasma use the Astrophysical Plasma Emission Code, apec [^3] APEC link https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node134.html code (Smith et al. 2001) with atomic data from atomdb v3.0.9 (Foster et al. 2012). This model uses the plasma temperature of the gas cells (in keV), the redshift $z$ and metallicity; Shreeram et al. (2024) assume a constant metallicity of $0.3\ Z_{\odot}$ for the generation of X-ray events. The X-ray events use the solar abundance values from Anders & Grevesse (1989). The events are generated by assuming a telescope with an energy-independent collecting area of $1000$ cm2 and an exposure time of $1000$ ks. The photon-list is generated in the observed frame of the X-ray emitting gas cells and is corrected to rest frame energies. Finally, the photons generated by the gas cells are projected onto the sky.
@@
We obtain X-ray radial surface brightness profiles in the $0.5-2.0$ keV band for all galaxies. Given that subfind provides us with an accurate classification of galaxies into centrals and satellites, we distinguish the hot gas component into X-ray emissions around centrals and satellites. For central galaxies, the profiles represent the hot gas emission around them; however, for satellite galaxies, the profiles probe the hot gas emission of the more massive host halo in the vicinity.
@@
We convolve the individual X-ray surface brightness profiles from LC-TNG300-based mock galaxies catalogues with the eROSITA PSF (Merloni et al. 2024). The PSF convolved mean X-ray surface brightness profile from hot gas, ${\mathbf{S}}_{\rm X,\ hot\ gas}[r]$, is expressed as follows.
$\centering\mathbf{S}_{\rm X,\ hot\ gas}(r)=f_{\rm cen}\ {S}_{\rm X,\ cen}+\mathcal{N}_{\rm sat}\times f_{\rm sat}\ {S}_{\rm X,\ sat},$ (3)
@@
where ${S}_{\rm X,\ cen}$ is the TNG-based prediction for the hot gas around central galaxies and ${S}_{\rm X,\ sat}$ corresponds to the hot gas around satellites. After matching LC-TNG300 with Fullphot in stellar mass and redshift, the mock galaxy catalogues fix the fraction of centrals, $f_{\rm cen}$, and satellites, $f_{\rm sat}$. $\mathcal{N}_{\rm sat}$ is the factor by which the mock prediction from ${S}_{\rm X,\ sat}$ is rescaled to match the observations, thereby renormalizes the ${S}_{\rm X,\ sat}$ of the TNG300-based prediction; $\mathcal{N}_{\rm sat}$ is the only free parameter in the hot gas emission component.
@@
The motivation behind introducing the renormalization parameter, $\mathcal{N}_{\rm sat}$, for fitting the forward model prediction for ${S}_{\rm X,\ sat}$ with observations are as follows. The TNG-based prediction for ${S}_{\rm X,\ sat}$ from the mock catalogues (for a given $f_{\rm sat}$) is 57$\sim$ 5-7$\times$ 5 - 7 × brighter than the Fullphot stack. Shreeram et al. (2024) find that the shape of the X-ray radial surface brightness profile from satellite galaxies (hosted by massive halos) is unaffected by $f_{\rm sat}$ in the galaxy sample. This is because the halo masses making up the average profile from satellite galaxies, whose $M_{\star}\in 10^{10.5-11}\ {\rm M}_{\odot}$, are dominated by host (central) halos with mean $M_{\rm 200m}\sim 10^{14}\ {\rm M}_{\odot}$. Therefore, by changing the normalization of the ${S}_{\rm X,\ sat}$, we effectively damp the normalization of the X-ray thermal gas contribution from the most massive clusters in the simulation. This is justified given that the hot gas fraction from TNG is overpredicted at halo masses above $M_{\rm 500c}\gtrsim 10^{13.5}\ {\rm M}_{\odot}$, as shown in Fig. 6 in Popesso et al. (2024a). This is also reflected in the $L_{\rm X}-M_{\rm 500c}$ relation shown in Zhang et al. (2024b) and Popesso et al. (2024b).
@@
We emphasize that the X-ray surface brightness profile prediction from central galaxies, ${S}_{\rm X,\ cen}$, which represents the CGM physics of interest in this work, is untouched. We predict multiple CGM profiles by changing the host halo mass distribution of the central galaxies and propagating it through our pipeline to generate mock galaxy catalogues for each halo distribution considered, as detailed in the following section. Note that the stellar mass and redshift distributions are the same for all three models.
@@
Table 1: Summary of the best-fit parameters (see Eq. 1-3) and derived quantities (luminosity values) obtained from fitting the three forward models from this work to the Fullphot X-ray surface brightness profile. For every model, we present the best-fit $\mathcal{N}_{\rm sat}$: renormalization of the ${S}_{\rm X,\ sat}$ (of the TNG300-based prediction), $f_{\rm sat}$: the fraction of satellites in the mock galaxy catalogue (see descriptions of the mock catalogues in Sec. 3.3), $\mathcal{N}_{\rm ps}$: the normalization of the point-source component (see Eq. 2), $L_{X,\rm\ PS}$: the X-ray luminosity obtained by integrating the point source component, $L_{X,\rm\ cenCGM}$: the X-ray luminosity obtained by integrating the central galaxies hot CGM component, $\chi^{2}_{\rm red}$: the reduced $\chi^{2}$ statistic for the model, obtained by using $21-2$ degrees of freedom.
@@
Model 2 $2.95_{-0.13}^{+0.14}$ 0.33 $1.1_{-0.2}^{+0.2}\times 10^{37}$ $1.195_{-0.042}^{+0.003}\times 10^{40}$ $2.76_{-0.19}^{+0.06}\times 10^{40}$ $1.32$
@@
3.3 Mock galaxy catalogues
We now use the LC-TNG300 galaxy catalogue to construct a mock galaxy sample for the LS DR9 Fullphot galaxies. We match every one of the $415,627$ galaxies in the Fullphot sample with a galaxy from LC-TNG300 in redshift and stellar mass. By construction, the simulated LC-TNG300 galaxies follow the same stellar mass and redshift distribution as the observational sample. The mock sample predicts the mean X-ray surface brightness profile for gas emitted around centrals and satellites.
In this work, we also test the impact of the underlying halo mass distribution on the CGM physics. Therefore, we additionally generate two other mock galaxy catalogues using LC-TNG300, matched in stellar mass and redshift, however, with different underlying halo mass distributions (see left panel of Fig. 1). Consequently, we also emulate the corresponding X-ray surface brightness profiles by varying the halo distributions (see right panel of Fig. 1). The differences between the three forward models are as follows:
@@ 3.3 Mock galaxy catalogues
By changing the underlying halo distributions for the fixed stellar mass bins, we are effectively changing the mean halo mass of our mock galaxy catalogue. We note that for model $3$, the mean halo mass is comparable with observational works that constrain the stellar-to-halo-mass relation in the low-redshift universe; for the mean stellar mass of our three mock samples ($M_{\star}=5.5\times 10^{10}\ {\rm M}_{\odot}$), observational works and empirical models predict between $M_{\rm halo}\sim 5\times 10^{11}\ {\rm M}_{\odot}$ (Taylor et al. 2020) to $1-3\times 10^{12}\ {\rm M}_{\odot}$ (Leauthaud et al. 2012; Coupon et al. 2015; Girelli et al. 2020; Behroozi et al. 2019). Given the large scatter in halo masses when selecting in fixed stellar mass bins (see e.g., Moster et al. 2020) and the stark dependence of the underlying halo distribution on the CGM profiles (right panel of Fig. 1), we are posteriorly constraining the properties of the halo distribution for the Fullphot dataset given our TNG-based model presented in this work.
Figure 2: Comparison of the mean point source (AGN and XRB) luminosities from our three forward models (crosses, based on the different halo distributions shown in Fig. 1) with the empirically allowed range of XRB and total point source luminosities, as shown by the grey hatched region and the green shaded region, respectively. We estimate the contribution due to XRB emission using the Aird et al. (2017) model. For estimating the AGN luminosity budget, ${L}_{X,\ \rm AGN}$, we use the Aird et al. (2013) model for the incidence rate distribution as a function of the $L_{X,\rm\ AGN}^{2-10\ \rm keV}$ keV. To covert the $2-10$ keV luminosity distribution in the $0.5-2$ keV band, we use an empirical obscuration model from Comparat et al. (2019). For more details, see the text of Sec. 4.1 and 4.2. This comparison favours model $3$, shown by the yellow cross, where the hot CGM component allows for a point source component with luminosity that agrees with empirical estimates from the low redshift universe.
@@ 4 Results and discussion
Figure 3: Decomposition of the X-ray stack of the galaxies in the photometric sample, Fullphot, into contributions from hot gas events (centrals and satellites hosted by more massive host halos) and point sources (AGN and XRB). The orange data points from Z24 are described with the model from this work (shown by the black solid line). The orange dashed-dotted line at $292$ kpc corresponds to the virial radius of the observational sample. The model is composed of the following: the hot CGM from central galaxies (yellow), the events around satellites probing the hot gas of their more massive host halos (green), and X-ray events from unresolved and resolved point-like sources comprising AGN and XRB (grey). The bottom panel shows the percentage deviation of the best-fit forward model from the data. The dashed-dotted lines show the $15\%$ level.
Figure 4: Posterior probability distributions of the renormalization factor of the ${S}_{\rm X,\ sat}$ profile: $\mathcal{N}_{\rm sat}$, and the normalization of the point source component: $\mathcal{N}_{\rm ps}$, which are obtained by fitting the forward-model $3$ from this work to the Fullphot data points from Z24 shown in Fig. 3. The vertical red lines in the diagonal plots correspond to the most likely value; the respective values are mentioned in the titles (refer to Tab. 1). The black dashed lines are the $68\%$ confidence interval of the marginalized distribution of the free parameters. The contour plot marks the most likely values with the red cross, and the contours correspond to the $68\%,95\%$ and $99.7\%$ confidence intervals.
4.1 Predicting the X-ray emission from XRB
The XRB emission, which is the X-ray emission from the binary component of stellar populations in normal galaxies, is divided into High-mass X-ray Binaries (HXRB) and Low-mass X-ray Binaries (LXRB); see review by Fabbiano (2006). The average XRB emission from a normal galaxy is characterized by scaling laws, where the former HXRB population scales with the recent star formation rate (SFR) in the galaxy (Grimm et al. 2003; Shtykovskiy & Gilfanov 2005; Mineo et al. 2012). In contrast, LXRB emission spans longer timescales, tracing the stellar mass of the galaxy (Gilfanov 2004; Boroson et al. 2011; Zhang et al. 2012; Lehmer et al. 2019). The total XRB emission from extragalactic objects is distributed on the scale of the stellar body; however, for an instrument with a $30$ arcsec PSF like eROSITA, it is unresolved and appears as a point source. Aird et al. (2017) and Lehmer et al. (2016) provide simple empirical recipes by parameterizing the total X-ray luminosity from XRB as a function of both the SFR and stellar mass, M, of the galaxy,

### v2

@@ 3.2 The hot gas component
In this work, we modeled the hot gas emission using the TNG300 hydrodynamical simulations (Pillepich et al. 2018; Marinacci et al. 2018; Naiman et al. 2018; Nelson et al. 2015; Springel et al. 2018); we used TNG300 to construct a lightcone and generate mock X-ray observations, as presented in Shreeram et al. (2025). Here, we summarize the most important features. We used the IllustrisTNG cosmological hydrodynamical simulation with the box of side length $302.6$ Mpc (Nelson et al. 2019, TNG300;) [^1] http://www.tng-project.org ; this box size allows us to map the hot CGM around MW-mass analogs embedded in the large-scale structure. TNG300 contains $2500^{3}$ dark matter particles, with a baryonic mass resolution of $1.1\times 10^{7}\rm\ M_{\odot}$ (resulting in $\gtrsim 10^{3}$ particles at MW-mass galaxies), a comoving value of the adaptive gas gravitational softening length for gas cells of $370$ comoving parsec (allowing us to resolve the X-ray gas from $\sim 5$ kpc from the halo center), gravitational softening of the collisionless component of $1.48$ kpc, and dark matter mass resolution of $5.9\times 10^{7}\ \rm M_{\odot}$. The TNG simulations adopt the Planck Collaboration XIII 2016 cosmological parameters. The TNG300 lightcone, LC-TNG300, was constructed with the box remap technique (Carlson and White 2010), using the LightGen code [^2] The code to generate lightcones from TNG is publicly available at https://github.com/SoumyaShreeram/LightGen/. LC-TNG300 spans across redshifts $0.03\lesssim z\lesssim 0.3$; this range is motivated by observations (e.g., Comparat et al. 2022; Chadayammuri et al. 2022; Zhang et al. 2024a). It goes out to $1231\rm\ cMpc$ along the x axis, subtending an area of $47.28\rm\ deg^{2}$ on the sky in the y-z plane. The physical properties of the distinct halos and subhalos within the TNG300 lightcone are obtained by the subfind algorithm (Springel et al. 2001; Dolag et al. 2009). subfind detects gravitationally bound substructures, equivalent to galaxies in observations, and also provides us with a classification of subhalos into centrals and satellites, where centrals are the most massive substructure within a distinct halo. For the MW-mass bin [^3] This paper defines the stellar mass used from TNG300 as the mass within twice the stellar half-mass radius., $\rm M_{\star}=10^{10.5-11}{\rm M}_{\odot}$, we have $5,109$ centrals and $2,719$ satellites, resulting in a total simulated galaxy catalog with $7,828$ galaxies (subhalos).
@@
The X-ray photons were simulated within the LC-TNG300 in the 0.52.00.5-2.0 keV intrinsic band with pyXsim (ZuHone and Hallman 2016), which is based on phox (Biffi et al. 2013, 2018b), by assuming an input emission model where the hot X-ray emitting gas is in collisional ionization equilibrium. The spectral model computations of hot plasma use the Astrophysical Plasma Emission Code, apec [^4] APEC link https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node134.html code (Smith et al. 2001) with atomic data from atomdb v3.0.9 (Foster et al. 2012). This model uses the plasma temperature of the gas cells (in keV), the redshift $z$, and metallicity; Shreeram et al. (2025) assume a constant metallicity of $0.3\ Z_{\odot}$ for the generation of X-ray events. The X-ray events use the solar abundance values from Anders and Grevesse (1989). The events are generated by assuming a telescope with an energy-independent collecting area of $1000$ cm2 and an exposure time of $1000$ ks. The photon list is generated in the observed frame of the X-ray emitting gas cells and is corrected to rest frame energies. Finally, the photons generated by the gas cells are projected onto the sky.
@@
We obtained X-ray radial surface brightness profiles in the 0.52.00.5-2.0 keV band for all galaxies. Given that subfind provides us with an accurate classification of galaxies into centrals and satellites, we distinguish the hot gas component into X-ray emissions around centrals and satellites. For central galaxies, the profiles represent the hot gas emission around them; however, for satellite galaxies, the profiles probe the hot gas emission of the more massive host halo in the vicinity.
@@
We convolved the individual X-ray surface brightness profiles from LC-TNG300-based mock galaxies catalogs with the eROSITA PSF (Merloni et al. 2024). The PSF convolved mean X-ray surface brightness profile from hot gas, ${\mathbf{S}}_{\rm X,\ hot\ gas}[r]$, is expressed as follows:
$\@add@centering\centering\mathbf{S}_{\rm X,\ hot\ gas}(r)=f_{\rm cen}\ {S}_{\rm X,\ cen}+\mathcal{N}_{\rm sat}\times f_{\rm sat}\ {S}_{\rm X,\ sat},$ (3)
@@
where ${S}_{\rm X,\ cen}$ is the TNG-based prediction for the hot gas around central galaxies and ${S}_{\rm X,\ sat}$ corresponds to the hot gas around satellites. After matching LC-TNG300 with Fullphot in stellar mass and redshift, the mock galaxy catalogs fix the fraction of centrals, $f_{\rm cen}$, and satellites, $f_{\rm sat}$. $\mathcal{N}_{\rm sat}$ is the factor by which the mock prediction from ${S}_{\rm X,\ sat}$ is rescaled to match the observations, thereby renormalizes the ${S}_{\rm X,\ sat}$ of the TNG300-based prediction; $\mathcal{N}_{\rm sat}$ is the only free parameter in the hot gas emission component.
@@
The motivation behind introducing the renormalization parameter, $\mathcal{N}_{\rm sat}$, for fitting the forward model prediction for ${S}_{\rm X,\ sat}$ with observations is as follows. The TNG-based prediction for ${S}_{\rm X,\ sat}$ from the mock catalogs (for a given $f_{\rm sat}$) is $\sim 5-7\times$ brighter than the Fullphot stack. Shreeram et al. (2025) find that the shape of the X-ray radial surface brightness profile from satellite galaxies (hosted by massive halos) is unaffected by $f_{\rm sat}$ in the galaxy sample. This is because the halo masses making up the average profile from satellite galaxies, whose $M_{\star}\in 10^{10.5-11}\ {\rm M}_{\odot}$, are dominated by host (central) halos with mean $M_{\rm 200m}\sim 10^{14}\ {\rm M}_{\odot}$. Therefore, by changing the normalization of the ${S}_{\rm X,\ sat}$, we effectively damp the normalization of the X-ray thermal gas contribution from the most massive clusters in the simulation. This is justified given that the hot gas fraction from TNG is overpredicted at halo masses above $M_{\rm 500c}\gtrsim 10^{13.5}\ {\rm M}_{\odot}$, as shown in Fig. 6 in Popesso et al. (2024a). This is also reflected in the $L_{\rm X}-M_{\rm 500c}$ relation shown in Zhang et al. (2024b) and Popesso et al. (2024b).
@@
We emphasize that the X-ray surface brightness profile prediction from central galaxies, ${S}_{\rm X,\ cen}$, which represents the CGM physics of interest in this work, is untouched. We predict multiple CGM profiles by changing the host halo mass distribution of the central galaxies and propagating it through our pipeline to generate mock galaxy catalogs for each halo distribution considered, as detailed in the following section. Note that the stellar mass and redshift distributions are the same for all three models.
We restricted our analysis to the MW-mass stellar bin due to both observational and modeling constraints. While comparisons across a broader stellar mass range are desirable, for lower-mass galaxies, there is currently no suitable X-ray stacking data available. For higher-mass galaxies, the existing stacked X-ray measurements (e.g., from Z24) extend to redshifts beyond the coverage of our lightcone, LC-TNG300, which is built from the TNG300 simulation. Extending the analysis to these massive galaxies would require constructing a new lightcone from a larger-volume simulation (e.g., FLAMINGO, MTNG), generating matched mock X-ray observations, and repeating the full forward-modeling pipeline introduced here, an effort that warrants a future study. In contrast, the MW-mass bin provides both observational data within the redshift range of LC-TNG300, making it the optimal case for robust model–data comparison and the hot CGM signal retrieval.
@@
Table 1: Summary of the best-fit parameters (see Eq. 1-3) and derived quantities (luminosity values) obtained from fitting the three forward models from this work to the Fullphot X-ray surface brightness profile. [^5] For every model, we present the best-fit $\mathcal{N}_{\rm sat}$: renormalization of the ${S}_{\rm X,\ sat}$ (of the TNG300-based prediction), $f_{\rm sat}$: the fraction of satellites in the mock galaxy catalog (see descriptions of the mock catalogs in Sec. 3.3), $\mathcal{N}_{\rm ps}$: the normalization of the point-source component (see Eq. 2), $L_{X,\rm\ PS}$: the X-ray luminosity obtained by integrating the point source component, $L_{X,\rm\ cenCGM}$: the X-ray luminosity obtained by integrating the central galaxies hot CGM component, $\chi^{2}_{\rm red}$: the reduced $\chi^{2}$ statistic for the model, obtained by using 21221-2 degrees of freedom.
@@
Model 2 $2.95_{-0.13}^{+0.14}$ 0.33 $1_{-0.2}^{+0.2}\times 10^{37} 1.1$$195_{-0.042}^{+0.003}\times 10^{40}$ $2.76_{-0.19}^{+0.06}\times 10^{40}$ $1.32$
@@
3.3 Mock galaxy catalogs
We next used the LC-TNG300 galaxy catalog to construct a mock galaxy sample for the LS DR9 Fullphot galaxies. We matched every one of the $415,627$ galaxies in the Fullphot sample with a galaxy from LC-TNG300 in redshift and stellar mass. By construction, the simulated LC-TNG300 galaxies follow the same stellar mass and redshift distribution as the observational sample. The mock sample predicts the mean X-ray surface brightness profile for gas emitted around centrals and satellites.
In this work, we also tested the impact of the underlying halo mass distribution on the CGM physics. Therefore, we additionally generated two other mock galaxy catalogs using LC-TNG300, matched in stellar mass and redshift; however, with different underlying halo mass distributions (see left panel of Fig. 1). Consequently, we also emulated the corresponding X-ray surface brightness profiles by varying the halo distributions (see right panel of Fig. 1). The differences between the three forward models are as follows:
@@ 3.3 Mock galaxy catalogs
[…]
Figure 2: Comparison of the mean point source (AGN and XRB) luminosities from our three forward models (crosses, based on the different halo distributions shown in Fig. 1) with the empirically allowed range of XRB and total point source luminosities, as shown by the hatched gray region and the shaded green region, respectively. We estimate the contribution due to XRB emission using the Aird et al. (2017) model. For estimating the AGN luminosity budget, ${L}_{X,\ \rm AGN}$, we used the Aird et al. (2013) model for the incidence rate distribution as a function of the $L_{X,\rm\ AGN}^{2-10\ \rm keV}$ keV. To convert the 2102-10 keV luminosity distribution in the 0.520.5-2 keV band, we used an empirical obscuration model from Comparat et al. (2019). For more details, see the text of Sec. 4.1 and 4.2. This comparison favors model $3$, shown by the yellow cross, where the hot CGM component allows for a point source component with luminosity that agrees with empirical estimates from the low-redshift Universe.
@@ 4 Results and discussion
4.1 Predicting the X-ray emission from XRBs
The XRB emission, which is the X-ray emission from the binary component of stellar populations in normal galaxies, is divided into high-mass X-ray binaries (HXRBs) and low-mass X-ray binaries (LXRBs); see review by Fabbiano (2006). The average XRB emission from a normal galaxy is characterized by scaling laws, where the former HXRB population scales with the recent star formation rate (SFR) in the galaxy (Grimm et al. 2003; Shtykovskiy and Gilfanov 2005; Mineo et al. 2012). In contrast, LXRB emission spans longer timescales, tracing the stellar mass of the galaxy (Gilfanov 2004; Boroson et al. 2011; Zhang et al. 2012; Lehmer et al. 2019). The total XRB emission from extragalactic objects is distributed on the scale of the stellar body; however, for an instrument with a $30$ arcsec PSF like eROSITA, it is unresolved and appears as a point source. Aird et al. (2017) and Lehmer et al. (2016) provide simple empirical recipes by parameterizing the total X-ray luminosity from XRBs as a function of both the SFR and stellar mass, M, of the galaxy,